5 Consider the investment project with the following net cash flows: Year Net Cash Flow 0 $1,500 SX $650 SX What would be the value of X if the projects IRR is known to be 10%? The annual income from a rented house is $24,000. The annual expenses are $6000. If the house can be sold for $245,000 at the end of 10 years, how much could you afford to pay for it now, if you considered 900 to be a suitable interest rate? 6 If you purchase a stock for $1500 now, what is the rate of return on your investment if the stock is worth $30037.50 at the end of 3 years? Find: i* 7 8. W hat is the effective annual rate of 20 percent compounded continuously? 9. You are borrowing money today at 8.48 percent, compounded annually. You will repay the principal plus all the interest in one lump sum of S12,800 two years from today. How much are you borrowing? 10 John loaned a friend $10,000 at 12% interest, compounded annually. The loan will be paid in four equal end-of-year payments. John expects the inflation rate to be 2%. After taking inflation into account, what rate of return is John receiving on the loan?
Answer #1

5.

NPV=0

-1500+X/(1+10%)^1+650/(1+10%)^2+X/(1+10%)^3=0

X*(1/(1+10%)^1+1/(1+10%)^3)=962.81

X=962.81/(1/(1+10%)^1+1/(1+10%)^3)

=579.86 or 580

the above is answer..

we do only one question based on Chegg rule